### v1

- STFT
- short-time Fourier transform
- RIR
- room impulse response
- iSTFT
- inverse short-time Fourier transform
- DoA
- direction of arrival
- IRM
- ideal ratio mask
- MVDR
- minimum-variance distortionless response
- NN
- neural network
- DNN
- deep neural network
- GCC-PHAT
- generalized cross-correlation phase transform
- VAD
- voice activity detection
- TdoA
- time difference of arrival
- SSL
- sound source localization
- SRP-PHAT
- steered response power with phase transform
- SLC
- spatial likelihood coding
- SSLC
- spatio-spectral likelihood coding
- MSE
- mean squared error
- MAE
- mean absolute error
- TCN
- time-convolutional network
- SI-SDR
- scale-invariant signal-to-distortion ratio
- PESQ
- perceptual evaluation of speech quality
- cACGMM
- complex angular central gaussian mixture model
- WER
- word error rate
- PIT
- permutation-invariant training
- LBT
- location-based training
- SSM
- spatial mixture model
- ASR
- automatic speech recognition
- DS
- direct separation
- DC
- deep clustering
- SCC
- spatial classification coding
- SSCC
- spatio-spectral classification coding
- LBOC
- location-based output channels
- LBOL
- location-based output likelihood
- BCE
- binary cross-entropy
- SRC
- spatial regression coding
- SSRC
- spatio-spectral regression coding
- JNF
- joint non-linear filter
- HAC
- hierachical agglomerative clustering
- SBC
- spatial binary coding
- SLC
- spatial likelihood coding
- MW-SBC
- mask-weighted spatial binary coding
- MW-SLC
- mask-weighted spatial likelihood coding
- NN
- neural network
- DNN
- deep neural network
- FB-MEst
- full-band mask estimator
- MC-CRUSE
- multi-channel convolutional recurrent U-net architecture for speech enhancement
- ESTOI
- extended short-time objective intelligibility
- SNR
- signal-to-noise ratio
Mask-Weighted Spatial Likelihood Coding for Speaker-Independent Joint Localization and Mask Estimation
Abstract
Due to their robustness and flexibility, neural-driven beamformers are a popular choice for speech separation in challenging environments with a varying amount of simultaneous speakers alongside noise and reverberation. Time-frequency masks and relative directions of the speakers regarding a fixed spatial grid can be used to estimate the beamformer’s parameters. To some degree, speaker-independence is achieved by ensuring a greater amount of spatial partitions than speech sources. In this work, we analyze how to encode both mask and positioning into such a grid to enable joint estimation of both quantities. We propose mask-weighted spatial likelihood coding and show that it achieves considerable performance in both tasks compared to baseline encodings optimized for either localization or mask estimation. In the same setup, we demonstrate superiority for joint estimation of both quantities. Conclusively, we propose a universal approach which can replace an upstream sound source localization system solely by adapting the training framework, making it highly relevant in performance-critical scenarios.
Index Terms— Multi-channel, speech separation, sound source localization, mask estimation, speaker-independent
1 Introduction
Given an audio recording containing multiple speakers, the field of speech separation deals with the simultaneous disentanglement of all speech sources into individual audio streams. While neural approaches have gained great proficiency in this task over the recent years, highly dynamic environments such as the so-called cocktail-party scenario [1] pose a great challenge until today. On top of noisy and reverberated speech signals, the varying amount of simultaneous speakers makes this problem especially demanding. In case recordings from a microphone array are available, the embedded spatial information can be leveraged to simplify the separation process. Especially the non-linear integration of spatial correlations and temporal-spectral patterns between the multi-channel signals from the array leads to powerful and computationally efficient neural network (NN) architectures [2, 3, 4, 5].
While popular concepts like permutation-invariant training (PIT) [6] or its multi-channel extension location-based training (LBT) [7] provide an efficient training framework, they are conceptually limited to a predetermined upper bound on simultaneous speech sources. This dependence arises from the fixed output size, which is deeply rooted in the architecture of the NN. While iterative separation approaches can circumvent this issue by extracting each speaker at a time in a spatially guided fashion [2, 4, 8, 9, 10, 11], the amount of subsequent NN executions in scenarios with many participants limits their applicability [12]. Neural sound source localization (SSL) systems [13] tackle the closely related problem of speaker-independent localization by assuming that all speakers are uniquely identifiable through their relative positioning towards the microphone array. By partitioning the recording environment into a fixed spatial grid, such approaches conduct either a binary or probabilistic speaker activity estimation for each individual region [14, 15]. Based on the success in SSL, this method has been adapted for speech separation. Instead of an activity indication, mask-weighted spatial binary coding (MW-SBC) encodes a time-frequency mask into each partition containing a speaker [16, 17, 18, 19]. However, in case of a very fine grid and sparse speech masks, there arises a significant label imbalance leading to an ill-conditioned optimization problem during training with regression-based loss functions [16, 18]. To avoid this issue, an alternative is to only utilize the output channels corresponding to the positions of the speakers. While this significantly improves the conditioning, it alleviates the need for the NN to conduct precise localization, thus a separate SSL system becomes necessary [18].
This work proposes a mask encoding which reduces the impact of label imbalance while at the same time retaining localization capabilities. By spectrally enriching the spatial likelihood coding (SLC) introduced in [15], we present mask-weighted spatial likelihood coding (MW-SLC) as alternative to MW-SBC and prove that it leads to a superior conditioning for a high spatial resolution. Finally, we show that MW-SLC dominates in joint localization and mask estimation while retaining considerable performance in both tasks individually.
2 Beamformer-guided Speech Separation
Let the speech mixture contain microphone recordings of stationary speech sources in a reverberant but noiseless environment. The signal propagation between -th source and all microphones shall be represented by the room impulse response (RIR) . In the short-time Fourier transform (STFT) domain at time frame and frequency bin , the acoustic scenario is resembled by
Due to its robustness and straightforward parameterization, we choose the minimum-variance distortionless response (MVDR) beamformer [20] for speech separation. Assuming knowledge about the microphone array geometry, it can be specified through the steering vector and covariance matrix . Dropping indices for visual clarity, the separated speech signals are given by
The steering vector is computed from the associated direction of arrival (DoA) assuming anechoic propagation between the microphones [20]. The second unknown, the spatial covariance matrix of the interference w.r.t. each speaker, is estimated by
in an interval of T frames with time-frequency masks . For this purpose we utilize ideal ratio masks [21, 22], which we threshold by to account for time-frequency bins without speech
Consequently both DoAs and IRMs are required for beamforming. Joint estimation of these quantities is the focus of this work.
3 Mask-Weighted Spatial Coding
3.1 Mask-Weighted Spatial Binary Coding
In this section we revisit mask-weighted spatial binary coding (MW-SBC) and analyze why regression-based loss functions lead to an ill-conditioned training [16, 18, 23]. MW-SBC divides the space around the microphone array into segments and assigns time-frequency masks to the directions where the speakers are present. The remaining segments are filled with zeros, so that the resulting encoding can be expressed as
with denoting the Kronecker delta [24], a binary indicator function which is non-zero only if equals . This can be seen as spectral extension of the binary DoA encoding [2, 25, 26, 27] which we will denote by spatial binary coding (SBC). In case a high spatial resolution is desired, the number of segments is much greater than the number of speakers . Consequently MW-SBC suffers from a significant label imbalance. Throughout the training of a NN, the resulting vast amount of zeros has a significant impact on the gradient during backpropagation. Especially regression-based loss functions, which are known to be robust against outliers, are prone to converge to poor, quasi-stationary solutions for the estimate of , such as [23]. Due to being a famous representative of regression-based loss functions, we will examine the mean squared error (MSE) more closely. Omitting the average about time and frequency bins, the MSE between and is defined by
| Specifically, we are interested in the gradient regarding , | ||||
| as it weights all differentiations w.r.t. the NN parameters due to the chain rule. Its norm regarding the spatial dimension yields | ||||
with representing the spatial vectorization of . Examining the norm at and inserting (5) gives
| Since the IRMs (4) are bounded by 1, the expression approaches zero in the limiting case of infinitely many partitions | ||||
Assuming the remaining differential chain is bounded, the sub-multiplicative property of norms [28] leads to a vanishing of the whole gradient w.r.t. the NN parameters. In other words, with a very large amount of segments , is correct in almost all cases. The remaining non-zero partitions containing IRMs only have negligible influence, which leads to a plateau in training. To avoid this problem, [16, 18] solely utilize the masks of active speakers during loss computation. While this removes the majority of zeros, it also decouples precise localization from the training objectives.
3.2 Proposed Mask-Weighted Spatial Likelihood Coding
In this subsection we present our proposed method to overcome the vanishing gradient problem of MW-SBC [18] and at the same time extend it towards joint localization and mask estimation. Our approach is inspired by the spatial likelihood coding (SLC) introduced in [15] within the context of SSL, which has gained a lot of popularity in recent SSL systems [8, 29, 30, 31]. The idea of SLC is to replace SBC by Gaussian curves centered at the DoAs. We propose to enrich this spatial likelihood with the spectral information from the IRMs by weighting the Gaussians accordingly. The resulting spatio-spectral coding, which we denote as mask-weighted spatial likelihood coding (MW-SLC), is defined as
with controlling the width of the Gaussians. The angular distance represents the wrapped mean absolute error (MAE) [32], thus taking the circularity of the DoAs into account. Similar to [15], we accumulate the individual curves with a maximum operation. However, in our case, due to weighting with the IRMs, the peaks of the underlying curves are no longer equally prominent. Especially for small angular distances and frequency bins dominated by a single speaker, adjacent masks can influence each other in our coding. While a chosen in correspondence to a minimum spatial gap between the speakers can alleviate this effect, it is a conceptual limitation MW-SBC does not suffer. To enable further analysis, we approximate the non-linear maximum operator in (11) with a sum
which is valid under the assumption of sharp and remote Gaussians. Based on this approximation, we conduct the same investigation for the MSE gradient norm at as for MW-SBC. Combining (8) with (11) at yields
| We expand this expression with the constant | ||||
with representing the angular space around the array in which the speakers can be located, e.g. in case of unconstrained speaker positioning . Considering the limiting case of a very high spatial resolution, thus the amount of segments approaches infinity, the partition size transitions to the differential . Together with the sum over the infinite amount of segments, the limit converges to the integral expression
Since the Gaussian function is strictly positive, the integral also evaluates to a positive non-zero value. Therefore, in contrast to MW-SBC in (10), the gradient’s norm of our proposed MW-SLC depends on the IRMs and is non-zero in speech presence. The main difference between both approaches lies in the continuous nature of the Gaussian encoding, which increases the spatial prominence with a rising amount of partitions , while it stays constant for the Kronecker delta function used in (5). To find a closed-form solution, we assume an unconstrained speaker positioning. By choosing the lower (equals upper) limit of integration centered between two DoAs, the circularity of the angular distance becomes negligible due to the assumption of remote and sharp Gaussians. In the same manner, the integration can be extended to the whole real number line, yielding |
The norm’s proportionality to underlines the strong connection between problem conditioning and width of the Gaussians. While an increased broadness improves training, it also puts less weight on accurate localization and possibly results in a reduced mask quality.
3.3 Joint DoA and Mask Extraction
Figure 1 summarizes the process of predicting with a deep neural network (DNN) and its application to both mask and DoA estimation. The latter is indicated on the left, where a peak-search w.r.t. is employed, the average of regarding all frequency bins ,
returning all angles at which is above a threshold and prominent in a neighborhood . Due to the frame-wise estimation, only the DoAs of the locally active speakers are found. This is represented by the estimated binary speech activity indicator . Therefore, to identify all DoAs in the recording, a post-processing step e.g. in form of clustering has to be applied. Finally, as indicated on the right-hand side of Fig. 1, the corresponding masks are extracted by slicing the estimated coding according to the DoA estimates, which we denote as sampling
| Model | Estimation | Localization | Separation | |||||||
| ID | Encoding | Architecture | Loss | DoA | IRM | MAE (∘) | Precision (%) | Recall (%) | SI-SDR (dB) | ESTOI (%) |
| (1) | - | Oracle | - | ✗ | ✗ | - / - | - / - | - / - | 6.60 / 7.33 | 69.2 / 57.9 |
| (2) | SBC | CNN/LSTM[25] | BCE | ✓ | ✗ | 3.33 / 6.82 | 81.4 / 84.7 | 94.7 / 75.2 | - / - | - / - |
| (3) | SLC | CNN/LSTM[25] | MSE | ✓ | ✗ | 0.63 / 2.36 | 93.3 / 94.0 | 95.2 / 83.3 | - / - | - / - |
| (4) | MW-SBC | FB-MEst[18] | MSE† | ✗ | ✓ | - / - | - / - | - / - | 5.26 / 6.38 | 63.6 / 54.5 |
| (5) | MW-SBC | MC-CRUSE[16] | MSE† | ✗ | ✓ | - / - | - / - | - / - | 5.85 / 6.53 | 65.5 / 55.0 |
| (6) | MW-SBC | FB-MEst[18] | MSE† | ✓ | ✓ | 20.74 / 28.53 | 52.0 / 39.3 | 60.3 / 55.5 | 3.08 / -0.25 | 60.8 / 48.5 |
| (7) | MW-SLC (ours) | FB-MEst | MSE | ✓ | ✓ | 1.26 / 3.05 | 89.4 / 90.1 | 94.5 / 85.4 | 4.99 / 5.92 | 62.8 / 53.0 |
| (8) | MW-SBC | MC-CRUSE[16] | MSE† | ✓ | ✓ | 20.65 / 28.61 | 45.0 / 36.8 | 48.6 / 44.5 | 0.26 / -1.54 | 59.9 / 48.8 |
| (9) | MW-SLC (ours) | MC-CRUSE | MSE | ✓ | ✓ | 2.32 / 7.24 | 88.8 / 88.4 | 92.1 / 77.6 | 5.03 / 5.64 | 62.6 / 52.3 |
4 Experiments
4.1 Experimental Setup
Dataset For training and evaluation we utilize the publicly available multi-channel dataset MC-LibriMix [8]. Specifically, we employ the versions with a sampling rate of 16-kHz generated for two (MC-Libri2Mix) and three (MC-Libri3Mix) speakers. The spatialization was conducted with simulated RIRs for a linear 4-channel microphone array (details in [8]). Regarding the STFT, we use a square-root Hann window [33] of length 32 ms and 16 ms hop-size.
DoA and Mask Estimation The width of the Gaussian curves in both SLC and MW-SLC is set to 6 as in [8]. The same value is used for the neighborhood in (17) as proposed in [15]. The threshold is determined via an exhaustive search on the validation dataset of MC-Libri2Mix w.r.t. optimizing the score. Finally, we utilize hierachical agglomerative clustering (HAC) to first obtain the DoAs and then the IRMs for all identified speakers. In each step we use the average distance between separate clusters as linkage and set as merging-condition. The threshold for the IRM computation in (4) is set to -35 dB.
4.2 Network Architectures and Training Details
NN Architecture To evaluate our proposed encoding, we employ two NN architectures introduced for MW-SBC, namely FB-MEst [18] and MC-CRUSE [16]. As a reference, we also list the SSL model CNN/LSTM [25] due to its close relation to FB-MEst. To further improve comparability, we utilize the mixture normalized regarding the channel dimension as input features [16, 18], see Fig. 1, and equip all architectures with Sigmoid output activations.
Training All NNs are trained on MC-Libri2Mix with a batch-size of 5 for 100 epochs or until convergence, which we define as no improvement on the validation dataset for 10 consecutive epochs. MC-Libri3Mix is only used during evaluation. The initial learning rate is set to 0.001 and reduced by a factor of 0.63 every 10 epochs, thus, leading to a decimation approximately every 50 epochs.
5 Results
Table 1 displays the results regarding SSL and joint localization and mask estimation in terms of separation performance with an MVDR beamformer.
Localization As it is common procedure in SSL, we assess the performance both in terms of a known (MAE) and unknown (precision, recall) amount of speakers according to the recipe from [15]. From two to three speakers almost all methods demonstrate increased precision and decreased recall scores. This can be traced back to optimizing the peak-search for the two speaker case, as the threshold is set too high to ideally accommodate more speakers. Regarding the baseline CNN/LSTM architecture, employing SLC as output coding (3) outperforms SBC (2) by a large margin. Possibly due to their similar NN architectures, FB-MEst trained with MW-SLC achieves almost the same localization results in (7). On the other hand, MC-CRUSE, which is originally not affiliated with SSL (5), displays slightly inferior performance (9). As expected, both NNs trained with MW-SBC (6, 8) perform significantly worse than with our proposed encoding in (7, 9), which can be also seen from a comparison between the DoA estimates in Fig. 2 (c) and (e).
Separation To evaluate speaker-independent separation performance, we follow the recipe proposed in [34]. The general gap between separation with an oracle IRM (1) and an estimated mask by employing MW-SBC (4, 5) can be attributed to a low input signal-to-noise ratio (SNR) of -5.50 dB for two and -8.18 dB for three speakers. Furthermore, the considerable difference between FB-MEst (4) and MC-CRUSE (5) can be linked to the architecture of FB-MEst, which leads to masks lacking the fine harmonic structure of speech, see [18]. Since the explicit localization criterion is dropped during training, MW-SBC supplied with oracle DoAs (4, 5) represents a natural upper bound for our proposed MW-SLC (7, 9). Regarding FB-MEst (7), both objective and perceptive measures come very close to this bound (4), however, the performance gap for MC-CRUSE (9) is more pronounced. As seen in Fig. 2, MW-SLC (f) does not match the temporal-spectral detail MW-SBC (d) achieves. We propose that this shortcoming could be improved by skewing the MW-SLC training loss with an additional term purely for mask reconstruction in future work. For completeness, we have also listed the separation results using estimated DoAs from MW-SBC (6, 8), which are clearly outperformed by our proposed method (7, 9).
6 Conclusion
In this work we proposed mask-weighted spatial likelihood coding (MW-SLC) for speaker-independent joint localization and mask estimation. Backed by theoretical investigations, we evaluated our approach towards encodings optimized for either localization or mask estimation. We showed that our method achieves considerable performance in both tasks w.r.t. the baselines, although performing them conjointly with equal computational overhead. In the same setup we demonstrated unmatched dominance for joint estimation. Conclusively, we proposed a universal approach which replaces an upstream SSL system by simply adapting the training scheme, making it highly relevant in performance-critical scenarios.

### v2

@@
 Conclusively, we propose a universal approach which can replace an upstream sound source localization system solely by adapting the training framework, making it highly relevant in performance-critical scenarios.
-Index Terms— Multi-channel, speech separation, sound source localization, mask estimation, speaker-independent
-1 Introduction
+Index Terms:
+Multi-channel, speech separation, sound source localization, mask estimation, speaker-independentI Introduction
 Given an audio recording containing multiple speakers, the field of speech separation deals with the simultaneous disentanglement of all speech sources into individual audio streams.
@@
 Especially the non-linear integration of spatial correlations and temporal-spectral patterns between the multi-channel signals from the array leads to powerful and computationally efficient neural network (NN) architectures [2, 3, 4, 5].
-While popular concepts like permutation-invariant training (PIT) [6] or its multi-channel extension location-based training (LBT) [7] provide an efficient training framework, they are conceptually limited to a predetermined upper bound on simultaneous speech sources.
-This dependence arises from the fixed output size, which is deeply rooted in the architecture of the NN.
+While popular concepts like permutation-invariant training (PIT) [6] or its multi-channel extension location-based training (LBT) [7] provide an efficient training framework, they are conceptually limited to a predetermined maximum amount of simultaneous speech sources.
+This dependency arises from the fixed output size, which is deeply rooted in the architecture of the NN.
 While iterative separation approaches can circumvent this issue by extracting each speaker at a time in a spatially guided fashion [2, 4, 8, 9, 10, 11], the amount of subsequent NN executions in scenarios with many participants limits their applicability [12].
@@
 Instead of an activity indication, mask-weighted spatial binary coding (MW-SBC) encodes a time-frequency mask into each partition containing a speaker [16, 17, 18, 19].
-However, in case of a very fine grid and sparse speech masks, there arises a significant label imbalance leading to an ill-conditioned optimization problem during training with regression-based loss functions [16, 18].
+However, in the case of a very fine grid and sparse speech masks, there arises a significant label imbalance leading to an ill-conditioned optimization problem during training with regression-based loss functions [16, 18].
 To avoid this issue, an alternative is to only utilize the output channels corresponding to the positions of the speakers.
@@
 Finally, we show that MW-SLC dominates in joint localization and mask estimation while retaining considerable performance in both tasks individually.
-2 Beamformer-guided Speech Separation
+II Beamformer-guided Speech Separation
 Let the speech mixture contain microphone recordings of stationary speech sources in a reverberant but noiseless environment.
 The signal propagation between -th source and all microphones shall be represented by the room impulse response (RIR) .
-In the short-time Fourier transform (STFT) domain at time frame and frequency bin , the acoustic scenario is resembled by
+In the short-time Fourier transform (STFT) domain at time frame and frequency bin , the acoustic scenario is thus resembled by
 Due to its robustness and straightforward parameterization, we choose the minimum-variance distortionless response (MVDR) beamformer [20] for speech separation.
-Assuming knowledge about the microphone array geometry, it can be specified through the steering vector and covariance matrix .
+Assuming knowledge about the microphone array geometry, it can be specified through the steering vector and the covariance matrix .
 Dropping indices for visual clarity, the separated speech signals are given by
@@
 Consequently both DoAs and IRMs are required for beamforming. Joint estimation of these quantities is the focus of this work.
-3 Mask-Weighted Spatial Coding
-3.1 Mask-Weighted Spatial Binary Coding
+III Mask-Weighted Spatial Coding
+III-A Mask-Weighted Spatial Binary Coding
 In this section we revisit mask-weighted spatial binary coding (MW-SBC) and analyze why regression-based loss functions lead to an ill-conditioned training [16, 18, 23].
@@
 This can be seen as spectral extension of the binary DoA encoding [2, 25, 26, 27] which we will denote by spatial binary coding (SBC).
-In case a high spatial resolution is desired, the number of segments is much greater than the number of speakers .
-Consequently MW-SBC suffers from a significant label imbalance.
+In case a high spatial resolution is desired, the number of segments is much greater than the number of speakers , thus, the majority of regions are empty.
+Additionally, due to the sparsity of speech signals in the time-frequency domain, the bins of the remaining partitions containing speakers are also predominantly inactive, which is why MW-SBC suffers from a significant label imbalance.
 Throughout the training of a NN, the resulting vast amount of zeros has a significant impact on the gradient during backpropagation.
@@
 | Specifically, we are interested in the gradient regarding , | ||||
-| as it weights all differentiations w.r.t. the NN parameters due to the chain rule. Its norm regarding the spatial dimension yields | ||||
-with representing the spatial vectorization of . Examining the norm at and inserting (5) gives
+| as it weights all differentiations w.r.t. the NN parameters according to the chain rule. Its norm regarding the spatial dimension yields | ||||
+| with representing the spatial vectorization of . Examining the norm at and inserting (5) gives | ||||
 | Since the IRMs (4) are bounded by 1, the expression approaches zero in the limiting case of infinitely many partitions | ||||
@@
 In other words, with a very large amount of segments , is correct in almost all cases.
-The remaining non-zero partitions containing IRMs only have negligible influence, which leads to a plateau in training.
+The remaining non-zero partitions containing IRMs have only negligible influence, which leads to a plateau in training.
 To avoid this problem, [16, 18] solely utilize the masks of active speakers during loss computation.
 While this removes the majority of zeros, it also decouples precise localization from the training objectives.
-3.2 Proposed Mask-Weighted Spatial Likelihood Coding
+III-B Proposed Mask-Weighted Spatial Likelihood Coding
 In this subsection we present our proposed method to overcome the vanishing gradient problem of MW-SBC [18] and at the same time extend it towards joint localization and mask estimation.
@@
 The resulting spatio-spectral coding, which we denote as mask-weighted spatial likelihood coding (MW-SLC), is defined as
-with controlling the width of the Gaussians.
+with the standard deviation of the Gaussians.
 The angular distance represents the wrapped mean absolute error (MAE) [32], thus taking the circularity of the DoAs into account.
@@
 To enable further analysis, we approximate the non-linear maximum operator in (12) with a sum
-which is valid under the assumption of sharp and remote Gaussians.
+which is valid under the assumption of Gaussians with low standard deviations and large mean difference.
 Based on this approximation, we conduct the same investigation for the MSE gradient norm at as for MW-SBC.
-Combining (8) with (11) at yields
-| We expand this expression with the constant | ||||
+Combining (8) with (12) at gives
+We rewrite this expression by introducing the constant
 with representing the angular space around the array in which the speakers can be located, e.g.
-in case of unconstrained speaker positioning .
+in the case of unconstrained speaker positioning .
 Considering the limiting case of a very high spatial resolution, thus the amount of segments approaches infinity, the partition size transitions to the differential .
@@
 To find a closed-form solution, we assume an unconstrained speaker positioning.
-By choosing the lower (equals upper) limit of integration centered between two DoAs, the circularity of the angular distance becomes negligible due to the assumption of remote and sharp Gaussians.
+By choosing the lower (equals upper) limit of integration centered between two DoAs, the circularity of the angular distance becomes negligible due to the assumption of Gaussians with low standard deviations and large mean difference.
 In the same manner, the integration can be extended to the whole real number line, yielding |
-The norm’s proportionality to underlines the strong connection between problem conditioning and width of the Gaussians.
-While an increased broadness improves training, it also puts less weight on accurate localization and possibly results in a reduced mask quality.
-3.3 Joint DoA and Mask Extraction
+The norm’s proportionality to underlines the strong connection between problem conditioning and standard deviation of the Gaussians.
+While an increased improves training, it also puts less weight on accurate localization and possibly results in a reduced mask quality.
+III-C Joint DoA and Mask Extraction
 Figure 1 summarizes the process of predicting with a deep neural network (DNN) and its application to both mask and DoA estimation.
@@
 | (1) | - | Oracle | - | ✗ | ✗ | - / - | - / - | - / - | 6.60 / 7.33 | 69.2 / 57.9 |
-| (2) | SBC | CNN/LSTM[25] | BCE | ✓ | ✗ | 3.33 / 6.82 | 81.4 / 84.7 | 94.7 / 75.2 | - / - | - / - |
-| (3) | SLC | CNN/LSTM[25] | MSE | ✓ | ✗ | 0.63 / 2.36 | 93.3 / 94.0 | 95.2 / 83.3 | - / - | - / - |
+| (2) | SBC | CNN/LSTM[25] | BCE | ✓ | - | 3.33 / 6.82 | 81.4 / 84.7 | 94.7 / 75.2 | - / - | - / - |
+| (3) | SLC | CNN/LSTM[25] | MSE | ✓ | - | 0.63 / 2.36 | 93.3 / 94.0 | 95.2 / 83.3 | - / - | - / - |
 | (4) | MW-SBC | FB-MEst[18] | MSE† | ✗ | ✓ | - / - | - / - | - / - | 5.26 / 6.38 | 63.6 / 54.5 |
@@
 | (9) | MW-SLC (ours) | MC-CRUSE | MSE | ✓ | ✓ | 2.32 / 7.24 | 88.8 / 88.4 | 92.1 / 77.6 | 5.03 / 5.64 | 62.6 / 52.3 |
-4 Experiments
-4.1 Experimental Setup
+IV Experiments
+IV-A Experimental Setup
 Dataset For training and evaluation we utilize the publicly available multi-channel dataset MC-LibriMix [8].
 Specifically, we employ the versions with a sampling rate of 16-kHz generated for two (MC-Libri2Mix) and three (MC-Libri3Mix) speakers.
-The spatialization was conducted with simulated RIRs for a linear 4-channel microphone array (details in [8]).
+The spatialization was conducted with simulated RIRs for a linear 4-channel microphone array.
+The simulations are constrained to reverberation times between 200 ms and 600 ms and a minimum angular distance between adjacent speakers of 15 ∘ (more details in [8]).
 Regarding the STFT, we use a square-root Hann window [33] of length 32 ms and 16 ms hop-size.
-DoA and Mask Estimation The width of the Gaussian curves in both SLC and MW-SLC is set to 6 as in [8].
+DoA and Mask Estimation The standard deviation of the Gaussian curves in both SLC and MW-SLC is set to 6 ∘ as in [8].
 The same value is used for the neighborhood in (18) as proposed in [15].
@@
 The threshold for the IRM computation in (4) is set to -35 dB.
-4.2 Network Architectures and Training Details
+IV-B Network Architectures and Training Details
 NN Architecture To evaluate our proposed encoding, we employ two NN architectures introduced for MW-SBC, namely FB-MEst [18] and MC-CRUSE [16].
@@
 The initial learning rate is set to 0.001 and reduced by a factor of 0.63 every 10 epochs, thus, leading to a decimation approximately every 50 epochs.
-5 Results
-Table 1 displays the results regarding SSL and joint localization and mask estimation in terms of separation performance with an MVDR beamformer.
+V Results
+Table I displays the results regarding SSL and joint localization and mask estimation in terms of separation performance with an MVDR beamformer.
+The loss computation is indicated by columns Encoding and Loss with SBC and SLC following [25] and [15] respectively.
 Localization As it is common procedure in SSL, we assess the performance both in terms of a known (MAE) and unknown (precision, recall) amount of speakers according to the recipe from [15].
@@
 For completeness, we have also listed the separation results using estimated DoAs from MW-SBC (6, 8), which are clearly outperformed by our proposed method (7, 9).
-6 Conclusion
+VI Conclusion
 In this work we proposed mask-weighted spatial likelihood coding (MW-SLC) for speaker-independent joint localization and mask estimation.